The top 5% of applicants (as measured by GRE scores) will receive scholarships to certain university. GRE scores ~ N(500, 100).

a) If an applicant who took GRE is selected at random, what is the probability that his/her score was more than $520 on that exam?

b) If an SRS of 100 applicants is selected, what is the probability that their average score on GRE is more than $520?

c) 99% of applicant will get between ____ and ____ points on GRE.

d)  How high does an applicant’s GRE score have to be to qualify for a scholarship?
Answer #1

solutions in l=500 G-100 S T = *- ? 2a) P(x - 520) = P123520-500) Too = P(230.2) 0.4207 <b> P/E3520) = P(23520-500) | Vio = P